# How connection probability shapes fluctuations of neural population dynamics

### v1

@@ II.3 Annealed network model
Figure 2: Schematic illustration of the annealed network and its approximation by a mean-connectivity network with additional noise. (a) In the annealed network model, Eq. (8), the random connectivity is re-sampled at each spike time resulting in an adjacency matrix $a_{ijk}$ that varies at each spike $k$. (b) The annealed model can be mapped to a mean-connectivity network with additional noise, Eq. (12). The network is all-to-all connected with an average connectivity $\langle a_{ij}\rangle$. Additionally, each neuron is subject to an independent noise that captures the random connectivity. (c) Example trajectories of the input potentials in the mean-connectivity network with additional noise (grey). The distribution $P$ of the input potentials follows a Gaussian distribution with a time-dependent mean $\bar{h}(t)$ and a time-dependent variance $\sigma^{2}(t)$. The second-order mean-field model (20) aims to describe the stochastic dynamics of these two mean-field variables.
@@
In the annealed approximation, the connections $a_{ij}$ are re-sampled at each spike occurring in the network. Thus, the time-varying adjacency matrix can be written as $[\hat{a}_{ijk}]$, where the additional index $k$ labels the $k$-th spike occurring in the network. Then, for a fixed $k$, the $N\times N$-adjacency matrix $[\hat{a}_{ij}]_{k}$ is still a random matrix with fixed in-degree, and therefore has independent rows but dependent elements $\hat{a}_{ijk}$ within a given row $i$. However, the dependence caused by the fixed-in-degree constraint is immaterial for the recurrent inputs in the annealed approximation. This is because only the $j_{k}$-th column of the matrix $[\hat{a}_{ij}]_{k}$, which has independent elements, matters for the transmission of the $k$-th spike, where $j_{k}$ is the index of the neuron that fired the $k$-th spike, and because of the re-sampling of the adjacency matrix. The latter implies that for different spike times, the columns are independent. From the perspective of the recurrent inputs, we can thus equally well re-sample the elements of the adjacency matrix as independent Bernoulli variables $a_{ijk}\sim\text{Ber}(p)$ with success probability $p=C/N$. Thus, our model, Eq. (1), changes in the annealed approximation to
$\tau dh_{i}(t)$ $=[-h_{i}+\mu(t)]dt+\frac{w}{C}dY_{i}(t-d),$ (8)
$\frac{dY_{i}}{dt}(t)$ $=\sum\limits_{j=1}^{N}\sum_{l}a_{ijk_{j,l}}\delta(t-t_{j,l})$ (9)
where the index $k_{j,l}$ delivers the location of the $l$-th spike time $t_{j,l}$ of neuron $j$ within the sequence of all spikes in the network. The quantity $Y_{i}(t)$ is the number of spikes received by neuron $i$ until time $t$ and, $dY_{i}/dt$ is the corresponding spike train. We refer to this model as the annealed network. We note that in simulations of the annealed network, it is not necessary to re-sample the entire adjacency matrix at each spike but only the column that corresponds to the neuron that fired the spike. In other words, whenever a neuron fires a spike, this spike is transmitted independently to postsynaptic neurons with probability $p=C/N$.
@@ III Derivation of the mesoscopic model
[…]
Here, $\Delta h_{i}(t)=h_{i}(t+\Delta t)-h_{i}(t)$ is the increment of the input potential and $\Delta Y_{i}(t)=Y_{i}(t+\Delta t)-Y_{i}(t)$ is the number of spikes received by neuron $i$ during the time step $\Delta t$. We now want to relate $\Delta Y_{i}(t)$ to the total number of spikes $\Delta Z(t)$ in the population during the same time step because the latter corresponds to the desired mesoscopic population activity via Eq. (3). The increment $\Delta Y_{i}(t)$ can be regarded as the result of a doubly stochastic process: On the one hand, given the total spike count $\Delta Z(t)$, each $\Delta Y_{i}(t)$ is the sum of $\Delta Z(t)$ independent Bernoulli variables according to Eq. (9) integrated over one time step. Hence, the increments $\Delta Y_{i}(t)$ are independent and binomially distributed with conditional mean and covariance
$\langle\Delta Y_{i}(t)|\Delta Z(t)\rangle=p\Delta Z(t),$
$\text{cov}(\Delta Y_{i}(t),\Delta Y_{j}(t)|\Delta Z(t))=p\left(1-p\right)\Delta Z(t)\delta_{ij}.$
If $N$ is large such that $p\Delta Z(t)\gg 1$, the number of synaptic input spikes $\Delta Y_{i}(t)$ for a given $\Delta Z(t)$ is approximately Gaussian distributed. Under this condition, we can therefore write
$\Delta Y_{i}(t)\approx p\Delta Z(t)+\sqrt{p\left(1-p\right)\Delta Z(t)}n_{i}(t),$ (11)
where $n_{i}(t)\sim\mathcal{N}(0,1)$ are independent standard normal random variables. The first term in Eq. (11) is proportional to the total number of spikes or the mesoscopic population activity and thus corresponds to the input in an effective fully connected network. This term is common to all neurons. In contrast, the second term is different among neurons and only arises when $p<1$. It thus captures the individual differences of the synaptic inputs in a non-fully connected (”diluted”) network.
@@
$\left\langle\Delta Y_{i}(t)\mid r(t)\right\rangle$ $=Cr(t)\Delta t$
$\text{cov}(\Delta Y_{i}(t),\Delta Y_{j}(t)\mid r(t))$ $=\left[p+(1-p)\delta_{ij}\right]Cr(t)\Delta t,$
where $\delta_{ij}$ denotes the Kronecker delta. The last equation shows that pairwise correlations of the synaptic input to different neurons are non-zero for $p>0$. These correlations arise from the first term in Eq. (11) representing the part of the recurrent input that is proportional to the population activity. The fluctuations of the population activity are common to all neurons, and thus yield a source of coherent noise.
@@
$\Delta\hat{Y}_{i}(t)=p\Delta Z(t)+\sqrt{C\left(1-p\right)r(t)\Delta t}n_{i}(t).$
@@
$\tau\frac{dh_{i}}{dt}=-h_{i}+\mu(t)+\frac{w}{N}\sum_{j=1}^{N}\dot{Z}_{j}(t-d)\\ +w\sqrt{\frac{1-p}{C}r(t-d)}\zeta_{i}(t).$ (12)
@@
Equation (12) represents an effective, fully connected network with rescaled synaptic efficacy $w/N=pJ$ and additional dynamic noise whose intensity is proportional to the variance $\sigma_{w}^{2}$ of the coupling strengths $w_{ij}$ (cf. Eq. (4)). Therefore, the dynamics is the same as the mean-connectivity network, Eq. (5), but with an additional noise term that captures the fluctuations caused by the random connectivity. This representation also reveals the distinct effects of finite network size ($N<\infty$) and random dilution ($p<1$) on the recurrent synaptic fluctuations. First, the stochastic input spikes from all neurons in the network contribute a common synaptic input current proportional to the population activity (third term on the right-hand side of Eq. (12)), which is stochastic for finite $N$ and thus exposes the coherent part of the synaptic fluctuations. More precisely, the population activity has conditional mean $r(t)$ (given all input potentials $h_{i}(t^{-})$, $i=1,\dotsc,N$) and finite-size fluctuations of order $N^{-1/2}$ because the spike trains $\dot{Z}_{i}(t)$ are conditionally Poisson (given $h_{i}(t^{-})$) with conditional mean $r(t)$.
In contrast, the last term of Eq. (12) contains a Gaussian white noise $\zeta_{i}(t)$ that is independent for each neuron, and thus represents the incoherent part of the synaptic fluctuations. Because the noise intensity is proportional to $\sigma_{w}^{2}=w^{2}(1-p)/p$, the noise term captures the effect of the quenched random connectivity ($\sigma_{w}^{2}>0$), and, specifically, the random dilution ($p<1$). It therefore vanishes in the limit of fully connected, homogeneous networks ($p\to 1$, $\sigma_{w}^{2}\to 0$), in which case the mean-connectivity network is recovered.
In the following, we assume a large network size $N$ such that a Gaussian approximation can be made: noting that $\Delta Z(t)$ is conditionally Poisson with mean and variance equal to $Nr(t)\Delta t$, the population activity can be approximated for large $N$ as
[…]
where $\eta(t)$ is a common Gaussian white noise with $\langle\eta(t)\eta(t^{\prime})\rangle=\delta(t-t^{\prime})$. In this form, the finite-size fluctuations of order $1/\sqrt{N}$ become explicit. We note that the Gaussian approximation of the population activity, Eq. (14), can in principle become negative. However, if $N$ is sufficiently large such that $Nr(t)\Delta t\gg 1$, the probability that such event happens is negligibly small.
@@ III.2 Second-order mesoscopic mean-field dynamics
The mean-connectivity network with additional noise derived in the last section is still an $N$-dimensional microscopic model and thus too complicated to analyze its dynamics. We now derive a mesoscopic mean-field model that is analytically tractable. To this end, we proceed in two steps: first, we consider the dynamics of the input potentials
@@
$g(t)$ $=w\sqrt{\frac{1-p}{C}r(t-d)}$
@@
with $\xi(t)=\sqrt{N}\left[\frac{1}{N}\sum_{i=1}^{N}\phi(h_{i}(t))-F\left(\bar{h}(t),\sigma^{2}(t)\right)\right]$. The rectification bracket $[\cdot]_{+}=\max(0,\cdot)$ could be inserted because $N^{-1}\sum_{i}\phi(h_{i})$ is always a non-negative quantity. Including the rectification is necessary for a Gaussian approximation of $\xi$ to enforce a non-negative population rate $r(t)$, albeit negative values are extremely rare events for biologically relevant network sizes $N$. For fixed input functions $f$ and $g$, the second term $F\left(\bar{h}(t),\sigma^{2}(t)\right)$ is deterministic, and hence, for large $N$, the variable $\xi(t)$ is an approximately Gaussian, colored noise with mean 00 and auto-covariance function $c_{\xi}^{f,g}(t,t^{\prime})=\text{cov}(\phi(h_{i}(t)),\phi(h_{i}(t^{\prime})))$ (again taken with fixed functions $f$ and $g$). Although there exist explicit expressions for the auto-covariance function of a nonlinear transformation of a Gaussian process with given auto-covariance function (see e.g. [65]), these expressions are in the form of infinite series or double integrals, and are not directly usable for the derivation of mesoscopic dynamics. Therefore, we follow a simpler ad hoc approach here: we assume that $\phi(h_{i}(t))$ has approximately the same temporal correlation structure as $h_{i}(t)$. Therefore, we make the heuristic approximation $c_{\xi}^{f,g}(t,t^{\prime})\approx\frac{c_{\xi}^{f,g}(t,t)}{\sigma^{2}}c_{h}^{f,g}(t,t^{\prime})$, where $c_{h}^{f,g}$ denotes the auto-correlation function of the process Eq. (15). Because $\xi(t)$ is Gaussian and $h_{i}(t)$ has an exponential auto-correlation function with correlation time $\tau$, we hence model the colored noise as an Ornstein-Uhlenbeck process of the form
@@
[…]
@@
We now proceed with the second step of the mean-field derivation, where we close the system in the mean-field variables $\bar{h}$, $\sigma^{2}$ and $\xi$ self-consistently. To this end, the functions $f$ and $g$ are related back to the stochastic mean-field rate $r(t)$ using Eq. (16). Inserting this equation into Eq. (17) we finally arrive at the mesoscopic mean-field dynamics
[…]
where $r(t)=[F\left(\bar{h}(t),\sigma^{2}(t)\right)+\xi(t)/\sqrt{N}]_{+}$. The population activity $A_{N}(t,\Delta t)$ associated with this mesoscopic dynamics is given by Eq. (14). Because we consider the population mean $\bar{h}(t)$ and population variance $\sigma^{2}(t)$, we call Eq. (20) the 2nd-order mean-field (MF) model.
The 2nd-order MF model reveals the effects of coherent finite-size fluctuations and incoherent fluctuations caused by the random disorder of the connectivity on the population dynamics. Firstly, finite-size noise appears in the equation for $\bar{h}$ as an explicit, multiplicative noise term, whose magnititude is proportional to $\sqrt{r/N}$. It thus increases with larger firing rates reflecting the Poisson noise in the underlying microscopic model, and it vanishes in the large-$N$ limit. Secondly, the incoherent fluctuations arising from the randomness of the connectivity are reflected by the equation for $\sigma^{2}$, which is essentially driven by the variance of the synaptic weights $\sigma_{w}^{2}=w^{2}(1-p)/p$. In turn, the variance $\sigma^{2}$ impacts the population rate $r$ through the averaged nonlinearity $F$, Eq. (18). However, the population average, Eq. (6), is not perfectly equal to $F$ for finite $N$ and $p<1$. The imperfect averaging leads to finite-size fluctuations $\xi(t)$ caused by both the finite network size and the random connectivity. This combined effect becomes manifest in our mean-field model by the noise strength proportional to $\sqrt{G/N}$ since $G$ vanishes for $\sigma^{2}=0$.
If we had neglected the dispersion of the input potentials due to the random connectivity, i.e. assuming $\sigma^{2}\equiv 0$, we would have arrived at a first-order MF model with only one equation for the mean $\bar{h}(t)$:

### v2

@@ II.3 Annealed network model
Figure 2: Schematic illustration of the annealed network and its approximation by a mean-connectivity network with additional noise. (a) In the annealed network model, Eq. (8), the random connectivity is re-sampled at each spike time resulting in an adjacency matrix $a_{ijk}$ that varies at each spike $k$. (b) The annealed model can be mapped to a mean-connectivity network with additional noise, Eq. (12). The network is all-to-all connected with an average connectivity $\langle a_{ij}\rangle$. Additionally, each neuron is subject to an independent noise that captures the random connectivity. (c) Example trajectories of the input potentials in the mean-connectivity network with additional noise (gray). The distribution $P$ of the input potentials follows a Gaussian distribution with a time-dependent mean $\bar{h}(t)$ and a time-dependent variance $\sigma^{2}(t)$. The second-order mean-field model (20) aims to describe the stochastic dynamics of these two mean-field variables.
@@
In the annealed approximation, the connections $a_{ij}$ are re-sampled at each spike occurring in the network. Thus, the time-varying adjacency matrix can be written as $\hat{a}_{ijk}]$, where the additional index $k$ labels the $k$-th spike occurring in the network. Then, for a fixed $k$, the $N\times N$-adjacency matrix $\hat{a}_{ij}]_{k}$ is still a random matrix with fixed in-degree, and therefore has independent rows but dependent elements $\hat{a}_{ijk}$ within a given row $i$. However, the dependence caused by the fixed-in-degree constraint is immaterial for the recurrent inputs in the annealed approximation. This is because only the $j_{k}$-th column of the matrix $\hat{a}_{ij}]_{k}$, which has independent elements, matters for the transmission of the $k$-th spike, where $j_{k}$ is the index of the neuron that fired the $k$-th spike, and because of the re-sampling of the adjacency matrix. The latter implies that for different spike times, the columns are independent. From the perspective of the recurrent inputs, we can thus equally well re-sample the elements of the adjacency matrix as independent Bernoulli variables $a_{ijk}\sim\text{Ber}(p)$ with success probability $p=C/N$. Thus, our model, Eq. (1), changes in the annealed approximation to
$\tau dh_{i}(t)$ $=[-h_{i}+\mu(t)]dt+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{N}}dY_{i}(t-d),$ (8)
$\frac{dY_{i}}{dt}(t)$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{w}{p}}\sum\limits_{j=1}^{N}\sum_{l}a_{ijk_{j,l}}\delta(t-t_{j,l})$ (9)
where the index $k_{j,l}$ delivers the location of the $l$-th spike time $t_{j,l}$ of neuron $j$ within the sequence of all spikes in the network. The quantity $Y_{i}(t)$ is the integrated synaptic input received by neuron $i$ until time $t$. We refer to this model as the annealed network (Table 1). We note that in simulations of the annealed network, it is not necessary to re-sample the entire adjacency matrix at each spike but only the column that corresponds to the neuron that fired the spike. In other words, whenever a neuron fires a spike, this spike is transmitted independently to postsynaptic neurons with probability $p=C/N$.
By generalizing from binary weights $w_{ij}=\frac{w}{p}a_{ij}$ to an arbitrary connectivity matrix $w_{ij}$, the annealed network can also be defined for such general connectivity through dynamic resampling, which produces time-dependent synaptic weights $w_{ij}(t)$. In the specific case of binary synapses, $w_{ij}(t_{j,l})=\frac{w}{p}a_{ijk_{j,l}}$ as in Eq. (9).
@@
Table 1: Summary of microscopic network models.
Network Equation Connectivity
Quenched (1) random, fixed in time,
$\left\langle w_{ij}\right\rangle=w$, $\text{var}(w_{ij})=\sigma_{w}^{2}$
Mean-connectivity (5) mean-matched all-to-all,
$\left\langle w_{ij}\right\rangle=w$, $\text{var}(w_{ij})=0$
Annealed (8), (9) random, dynamic
(resampling in time),
$\left\langle w_{ij}(t)\right\rangle=w$, $\text{var}(w_{ij}(t))=\sigma_{w}^{2}$
@@ III Derivation of the mesoscopic model
$\tau\Delta h_{i}(t)=\int_{t}^{t+\Delta t}[-h_{i}(s)+\mu(s)]\,ds+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{N}}\Delta Y_{i}(t-d).$ (10)
Here, $\Delta h_{i}(t)=h_{i}(t+\Delta t)-h_{i}(t)$ is the increment of the input potential, $\Delta Y_{i}(t)=\sum_{j=1}^{N}w_{ij}(t)\Delta Z_{j}(t-d)$ is the synaptic input received by neuron $i$ during the time step $\Delta t$ and $\Delta Z_{j}(t)=1$ if neuron $j$ has a spike during that time step and $\Delta Z_{j}(t)=0$ otherwise. In the annealed approximation, the time-varying coupling strengths $w_{ij}(t)$ are independent random variables (across time steps and synapses) with mean $w$ and variance $\sigma_{w}^{2}$. In our specific example, the random variables are binary taking the value $w/p$ with probability $p$ and 0 otherwise. However, the derivation does not hinge upon binary connectivity but applies more generally to any distribution of coupling strengths with finite variance $\sigma_{w}^{2}<\infty$.
The synaptic input $\Delta Y_{i}(t)$ can be regarded as the result of a doubly stochastic process: On the one hand, given the total spike count $\Delta Z(t)=\sum_{i=1}^{N}\Delta Z_{i}(t)$, each $\Delta Y_{i}(t)$ is the sum of $\Delta Z(t)$ independent random variables. Hence, the conditional mean and covariance are
$\langle\Delta Y_{i}(t)|\Delta Z(t)\rangle={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}w}\Delta Z(t),$
$\text{cov}(\Delta Y_{i}(t),\Delta Y_{j}(t)|\Delta Z(t))={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{w}^{2}}\Delta Z(t)\delta_{ij},$
where $\delta_{ij}$ denotes the Kronecker delta. If $N$ is large such that $p\Delta Z(t)\gg 1$, the synaptic input $\Delta Y_{i}(t)$ for a given $\Delta Z(t)$ is approximately Gaussian distributed. Under this condition, we can write
$\Delta Y_{i}(t)\approx{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}w}\Delta Z(t)+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{w}}\sqrt{\Delta Z(t)}n_{i}(t),$ (11)
where $n_{i}(t)\sim\mathcal{N}(0,1)$ are independent standard normal random variables. The first term in Eq. (11) is proportional to the total number of spikes or the mesoscopic population activity and thus corresponds to the input in an effective mean-connectivity network. This term is common to all neurons. In contrast, the second term is independent across neurons and only arises when $\sigma_{w}>0$. Specifically, it captures the individual differences of the synaptic inputs in a non-fully connected (“diluted”) network with $p<1$.
@@
$\left\langle\Delta Y_{i}(t)\mid r(t)\right\rangle$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}wN}r(t)\Delta t$
$\text{cov}(\Delta Y_{i}(t),\Delta Y_{j}(t)\mid r(t))$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(w^{2}+\sigma_{w}^{2}\delta_{ij}\right)N}r(t)\Delta t.$
The last equation shows that pairwise correlations of the synaptic input to different neurons are non-zero for finite $N$. These correlations arise from the common fluctuations of the first term in Eq. (11) representing the population activity.
@@
$\Delta\hat{Y}_{i}(t)={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}w}\Delta Z(t)+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{w}}\sqrt{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}N}r(t)\Delta t}n_{i}(t).$
@@
$\tau\frac{dh_{i}}{dt}=-h_{i}+\mu(t)+\frac{w}{N}\sum_{j=1}^{N}\dot{Z}_{j}(t-d)+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{w}\sqrt{\frac{r(t-d)}{N}}}\zeta_{i}(t).$ (12)
@@
Equation (12) represents an effective, fully connected network with rescaled synaptic efficacy $w/N=pJ$ and additional dynamic noise whose intensity is proportional to the variance $\sigma_{w}^{2}$ of the coupling strengths $w_{ij}$. Therefore, the dynamics is the same as the mean-connectivity network, Eq. (5), but with an additional noise term that captures the fluctuations caused by the random connectivity. This representation also reveals the distinct effects of finite network size ($N<\infty$) and random dilution ($p<1$) on the recurrent synaptic fluctuations. First, the stochastic input spikes from all neurons in the network contribute a common synaptic input current proportional to the population activity (third term on the right-hand side of Eq. (12)), which is stochastic for finite $N$ and thus exposes the coherent part of the synaptic fluctuations. More precisely, the population activity has conditional mean $r(t)$ (given all input potentials $h_{i}(t^{-})$, $i=1,\dotsc,N$) and finite-size fluctuations of order $N^{-1/2}$ because the spike trains $\dot{Z}_{i}(t)$ are conditionally Poisson (given $h_{i}(t^{-})$) with conditional mean $r(t)$.
In contrast, the last term of Eq. (12) contains a Gaussian white noise $\zeta_{i}(t)$ that is independent for each neuron, and thus represents the incoherent part of the synaptic fluctuations. This incoherent noise captures the effect of the random connectivity ($\sigma_{w}^{2}>0$), and, specifically, the random dilution ($p<1$). It therefore vanishes in the limit of fully connected, homogeneous networks ($p\to 1$, $\sigma_{w}^{2}\to 0$), in which case the mean-connectivity network is recovered.
In the following, we assume a large network size $N$ such that a diffusion approximation of the shot noise can be made: noting that $\Delta Z(t)$ is conditionally Poisson with mean and variance equal to $Nr(t)\Delta t$, the population activity can be approximated for large $N$ as
$A_{N}(t;\Delta t)\approx\frac{\int_{t}^{t+\Delta t}A(t^{\prime})\,dt^{\prime}}{\Delta t},\quad A(t)=r(t)+\sqrt{\frac{r(t)}{N}}\eta(t),$ (14)
where $\eta(t)$ is a common Gaussian white noise with $\langle\eta(t)\eta(t^{\prime})\rangle=\delta(t-t^{\prime})$. In this form, the finite-size fluctuations of order $1/\sqrt{N}$ become explicit. We note that the diffusion approximation of the population activity, Eq. (14), can in principle become negative. However, if $N$ is sufficiently large such that $Nr(t)\Delta t\gg 1$, the probability that such event happens is negligibly small.
@@ III.2 Second-order mesoscopic mean-field dynamics
All-to-all uniformly connected networks with independent noise, such as the mean-connectivity network with additional noise derived in the previous section, have been well studied. In particular, in the limit $N\to\infty$, the population dynamics can be described by a Fokker-Planck equation [70], while for finite $N$, the stationary statistics of the fluctuations has been calculated using linear response theory [71]. However, despite earlier theoretical approaches [29, 66], a derivation of stochastic population dynamics that captures both finite-size fluctuations and nonstationary dynamics remains challenging. In our setting, this derivation is facilitated by the absence of resets in the Poisson neuron model. To obtain a self-consistent description, we follow a standard two-step procedure: first, we consider the dynamics of the input potentials
@@
$g(t)$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{w}\sqrt{\frac{r(t-d)}{N}}}$
@@
with $\xi(t)=\sqrt{N}\left[\frac{1}{N}\sum_{i=1}^{N}\phi(h_{i}(t))-F\left(\bar{h}(t),\sigma^{2}(t)\right)\right]$. The rectification bracket $\cdot]_{+}=\max(0,\cdot)$ could be inserted because $N^{-1}\sum_{i}\phi(h_{i})$ is always a non-negative quantity. Including the rectification is necessary for a Gaussian approximation of $\xi$ to enforce a non-negative population rate $r(t)$, albeit negative values are extremely rare events for biologically relevant network sizes $N$. For fixed input functions $f$ and $g$, the second term $F\left(\bar{h}(t),\sigma^{2}(t)\right)$ is deterministic, and hence, for large $N$, the variable $\xi(t)$ is an approximately Gaussian, colored noise with mean 0 and auto-covariance function $c_{\xi}^{f,g}(t,t^{\prime})=\text{cov}(\phi(h_{i}(t)),\phi(h_{i}(t^{\prime})))$ (again taken with fixed functions $f$ and $g$). Although there exist explicit expressions for the auto-covariance function of a nonlinear transformation of a Gaussian process with given auto-covariance function (see e.g. [72]), these expressions are in the form of infinite series or double integrals, and are not directly usable for the derivation of mesoscopic dynamics. Therefore, we follow a simpler ad hoc approach here: we assume that $\phi(h_{i}(t))$ has approximately the same temporal correlation structure as $h_{i}(t)$. Therefore, we make the heuristic approximation $c_{\xi}^{f,g}(t,t^{\prime})\approx\frac{c_{\xi}^{f,g}(t,t)}{\sigma^{2}}c_{h}^{f,g}(t,t^{\prime})$, where $c_{h}^{f,g}$ denotes the auto-correlation function of the process Eq. (15). Because $\xi(t)$ is Gaussian and $h_{i}(t)$ has an exponential auto-correlation function with correlation time $\tau$, we hence model the colored noise as an Ornstein-Uhlenbeck process of the form
@@
$G\left(\bar{h},\sigma^{2}\right):=\int_{-\infty}^{\infty}\phi^{2}(h)g_{\bar{h},\sigma^{2}}(h)\,dh-F^{2}\left(\bar{h},\sigma^{2}\right).$
@@
We now proceed with the second step of the mean-field derivation, where we close the system in the mean-field variables $\bar{h}$, $\sigma^{2}$ and $\xi$ self-consistently. To this end, the functions $f$ and $g$ are related back to the stochastic mean-field rate $r(t)$ using Eq. (16). Inserting this equation into Eq. (17) we finally arrive at the mesoscopic mean-field (MF) dynamics
$\begin{aligned} \tau\frac{\mathrm{d}\bar{h}}{\mathrm{d}t}&=-\bar{h}+\mu(t)+w\left[r(t-d)+\sqrt{\frac{r(t-d)}{N}}\eta(t)\right]\\ \tau\frac{\mathrm{d}\sigma^{2}}{\mathrm{d}t}&=-2\sigma^{2}+\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{w}^{2}}}{\tau N}r(t-d)\\ \tau\frac{\mathrm{d}\xi}{\mathrm{d}t}&=-\xi+\sqrt{2\tau G\left(\bar{h}(t),\sigma^{2}(t)\right)}\zeta(t),\end{aligned}$ (20)
where $r(t)=[F\left(\bar{h}(t),\sigma^{2}(t)\right)+\xi(t)/\sqrt{N}]_{+}$. The population activity $A_{N}(t,\Delta t)$ associated with this mesoscopic dynamics is given by Eq. (14). Because we consider the population mean $\bar{h}(t)$ and population variance $\sigma^{2}(t)$, we call Eq. (20) the 2nd-order mean-field (MF2) dynamics.
The MF2 dynamics captures the effects of both finite $N$ and random connectivity. Firstly, the finite-size effect manifests itself in the multiplicative noise term proportional to $\sqrt{r/N}$ in the equation for $\bar{h}$. The finite-size noise increases with larger firing rates, reflecting the Poisson noise in the underlying microscopic model, and it vanishes in the large-$N$ limit. Secondly, the effect of random connectivity becomes manifest in a non-vanishing variance $\sigma^{2}$, which is driven by the term proportional to $\sigma_{w}^{2}/N$. Note that in our case of binary random connectivity, the variance of the synaptic weights is $\sigma_{w}^{2}=w^{2}(1-p)/p$ (cf. Eq. (4)). Thus, for full connectivity, $p=1$, the variance $\sigma^{2}$ vanishes. It also vanishes for $N\to\infty$ unless $C$ is kept constant (sparse limit) in which case $\sigma_{w}^{2}/N=w^{2}(1/C-1/N)$ remains of order $1$.
A non-vanishing variance $\sigma^{2}$ (in our case $p<1$), in turn, impacts the population rate $r$ through the averaged nonlinearity $F$, Eq. (18). However, the population average, Eq. (6), is not perfectly equal to $F$ for finite $N$ and $\sigma^{2}>0$. The imperfect averaging leads to finite-size fluctuations $\xi(t)$ – a third effect caused by the combination of both finite network size and random connectivity. This effect vanishes when $N\to\infty$ or $\sigma_{w}^{2}=0$ ($p=1$) because the noise strength is proportional to $\sqrt{G/N}$ and $G$ vanishes for $\sigma^{2}=0$.
If we had neglected the dispersion of the input potentials due to the random connectivity, i.e. assuming $\sigma^{2}\equiv 0$, we would have arrived at a first-order MF theory (MF1) with only one equation for the mean $\bar{h}(t)$: